At 101.4 oC, the Kp for the reaction:

1A+4B ↔ 2AB

is 1.25⋅10^0

What is Keq? Enter your answer in scientific notation.
Answer #1

T= 101.4 oC
= (101.4+273) K
= 374.4 K
Δ n = number of gaseous molecule in product - number of gaseous molecule in reactant
Δ n = -3

use:
Kp= Kc (RT)^Δn
1.25 = Kc *(0.08206*374.4)^(-3)
Kc = 3.625*10^4
Answer: 3.62*10^4